### v1

# Calib3D: Calibrating Model Preferences for Reliable 3D Scene Understanding

Abstract
Safety-critical 3D scene understanding tasks necessitate not only accurate but also confident predictions from 3D perception models. This study introduces Calib3D, a pioneering effort to benchmark and scrutinize the reliability of 3D scene understanding models from an uncertainty estimation viewpoint. We comprehensively evaluate 28 state-of-the-art models across 10 diverse 3D datasets, uncovering insightful phenomena that cope with both the aleatoric and epistemic uncertainties in 3D scene understanding. We discover that despite achieving impressive levels of accuracy, existing models frequently fail to provide reliable uncertainty estimates – a pitfall that critically undermines their applicability in safety-sensitive contexts. Through extensive analysis of key factors such as network capacity, LiDAR representations, rasterization resolutions, and 3D data augmentation techniques, we correlate these aspects directly with the model calibration efficacy. Furthermore, we introduce DeptS, a novel depth-aware scaling approach aimed at enhancing 3D model calibration. Extensive experiments across a wide range of configurations validate the superiority of our method. We hope this work could serve as a cornerstone for fostering reliable 3D scene understanding. Code and benchmark toolkits are publicly available. [^1] GitHub Repo: https://github.com/ldkong1205/Calib3D.
Keywords:
3D Semantic Segmentation Uncertainty Estimation Network Calibration 3D Perception Robustness Autonomous Driving
1 Introduction
The reliability of the perception system in real-world conditions is paramount. Safety-critical applications, such as autonomous driving and robot navigation, often rely on robust and accurate predictions from perception models [130, 98].
Albeit being the most popular choice, learning-based perception models are known to struggle in delivering uncertainty estimates [34] and suffer from over- or under-confidences [47]. In other words, prior models tend to be badly calibrated, which cannot satisfy the needs in real life [1, 79]. This is contradictory to the objective of safe operations of autonomous vehicles, where the perception models are anticipated to predict the presence or absence of obstacles in a fairly confident manner [97, 55, 116]. A similar rationale often applies to other safety-critical applications, such as surveillance [62, 77, 88], automated health care [74, 50, 73], remote sensing [94, 33], etc., where the trustworthiness plays the critical role.
Figure 1: Well-calibrated 3D scene understanding models are anticipated to deliver low uncertainties when predictions are accurate and high uncertainties when predictions are inaccurate. Existing 3D models [133] (UnCal) and prior calibration methods [36, 71] struggled to provide proper uncertainty estimates. Our proposed depth-aware scaling (DeptS) is capable of outputting accurate estimates, highlighting its potential for real-world usage. The plots shown are the point-wise expected calibration error (ECE) rates. The colormap goes from dark to light denoting low and high error rates, respectively.
Several studies have attempted to understand the reliability of image recognition models and observed insightful phenomenons [61, 71, 80, 84]. Guo et al. [36] presented one of the first benchmarks for network calibration, revealing the fact that modern neural networks are no longer well-calibrated. Subsequent works stemmed from similar motivations and drew similar conclusions for other mainstream image-based perception tasks, including object detection [62, 77, 52, 81, 88], depth estimation [53, 48, 87, 108], and image segmentation [24, 51, 110, 76, 9].
Motivation. Despite existing endeavors, the reliability of 3D scene understanding models in delivering uncertainties [^2] In this work, for the sake of clarity, the terms uncertainty and confidence are used interchangeably, i.e., uncertainty $=1~{}-$ confidence. remains unknown. 3D data, such as those acquired by LiDAR sensors and RGB-D cameras, are much sparser and often lack structures than images [11, 2, 114]. Calib3D emerges as a comprehensive study dedicated to scrutinizing and benchmarking the reliability of 3D scene understanding models through the lens of uncertainty estimation. Our exploration delves into both aleatoric and epistemic uncertainties, aiming to unravel the nuanced dynamics of 3D perception models and their calibration needs in real-world, safety-critical settings. The major focuses of this study are:
  • Aleatoric Uncertainty in 3D. We examine how intrinsic factors, such as sensor measurement noises [31] and point cloud density variations [55, 105], contribute to data uncertainty in 3D scene understanding. Such uncertainty cannot be reduced even with more data or improved models, necessitating efforts that effectively interpret and quantify this inherent variability. Calib3D encompasses a total of 10 popular 3D datasets, ranging from diverse sensor acquisitions, annotation protocols, and scene conditions. Our benchmark covers diverse 3D scenarios, including driving, off-road, and indoor scenes, as well as dynamic, synthetic, corrupted, and adverse weather conditions.
  • Epistemic Uncertainty in 3D. Different from the rather unified network structures in 2D [40, 25], 3D scene understanding models encompass a wider array of structures due to the complex nature of 3D data processing. Our investigation in Calib3D extends to the model uncertainty associated with the diverse 3D architectures, highlighting the importance of addressing knowledge gaps in model training and data representation. A total of 28 state-of-the-art models are compared and analyzed, shedding light on the future development of more reliable 3D scene understanding models.
Our investigation into the current landscape of 3D scene understanding models reveals a critical gap: while these models often achieve promising levels of accuracy, their calibration abilities –essential for trust in safety-critical applications – consistently fall short of the mark. This discrepancy, as visualized in Fig. 1, highlights the need for more sophisticated calibration methods that can align model confidence with their predictive accuracy. In-depth analysis of various factors, including network capacity, LiDAR data representations, rasterization resolutions, and 3D data augmentation techniques, has led us to identify key areas where calibration performance can be significantly improved.
To further enhance uncertainty estimation capabilities, we propose a depth-aware scaling method, dubbed DeptS. Our method is motivated by the observation that uncalibrated models tend to have low accuracy in the middle-to-far region of the ego-vehicle, while, in the meantime, posing severe over-confident predictions. This problem, which is directly correlated with 3D scene structural information, inevitably leads to high calibration errors. To tackle this challenge, we design a depth-correlated temperature to dynamically adjust the logits distribution based on the depth information, exhibiting strong generalizability in calibrating 3D perception models. As shown in Fig. 1, DeptS not only demonstrates appealing calibration performance over the uncalibrated model but also outperforms several calibration methods from existing literature [36, 61, 110, 71].
To encapsulate, this work is featured by the following seminar contributions:
  • To the best of our knowledge, Calib3D is the first benchmark targeted at probing the uncertainties of 3D perception models in real-world conditions.
  • Our benchmark encompasses a systematic study of 28 state-of-the-art models across 10 diverse 3D datasets, laying a solid foundation for the future research and development of reliable 3D scene understanding models.
  • We proposed DeptS, a straightforward yet effective depth-aware scaling method that better calibrates the uncertainty estimates for 3D models.
  • Extensive experimental evaluations across a wide range of 3D scenarios demonstrate our advantages, shedding light on a more reliable 3D scene understanding that extends well beyond the current state of the art.
2 Related Work
3D Scene Understanding. Holistic perception of 3D scenes lays the foundation of many real-world applications [11, 6]. Based on the choice of 3D representations [106], existing approaches can be categorized into range view [75, 131, 16, 118, 57, 54, 122], bird’s eye view [129, 14, 132], sparse voxel [18, 133, 101, 44, 43, 17], and point-based [45, 103, 128, 89] methods. Most recent works seek to combine the advantages of multiple representations together [65, 119, 67] or fuse other modalities (e.g., cameras, Radars, and IMU) with point clouds [134, 49, 68, 85, 120] for more accurate 3D perceptions. Although promising performance has been achieved on popular 3D scene understanding datasets, the reliability of existing 3D perception models in estimating the predictive uncertainty remains unknown.
Uncertainty Estimation. Quantifying uncertainties is crucial in real-world scenarios [4, 41]. For safety-critical applications like 3D scene understanding, it is especially critical to obtain reliable estimates of uncertainties [82]. In general, methods on the uncertainty estimation track can be categorized into various types, i.e., deterministic networks [72, 96], Bayesian methods [23, 104, 10, 42, 30], ensembles [63, 92, 37], and test-time augmentations [70, 5]. Recent studies pay attention to post-hoc approaches for calibrating uncertainties, which align with practical usages [34]. This work follows this line of research and extends efforts to 3D scene understanding, hoping to enlighten future works on this crucial topic.
Network Calibration. As one of the prevailing research topics, numerous calibration methods have been proposed to tackle different tasks, including image classification [36, 127, 61, 71, 80, 84, 60], semantic segmentation [24, 51, 110, 76, 9, 27], object detection [62, 77, 52, 81, 88], depth estimation [53, 48, 87, 108], remote sensing [94, 33], medical imaging [74, 50, 73], and so on. However, the calibration of 3D scene understanding models, which relates closely to real-world applications, is rather overlooked in the literature. Dreissig et al. [28] made a preliminary study of the uncertainties in SalsaNext [21] on SemanticKITTI [7]. As far as we know, Calib3D serves as the first systematic study of uncertainty estimation for 3D scene understanding. Our benchmark covers a total of 28 state-of-the-art models across 10 diverse 3D datasets, paving the way for the future development of more reliable and practical 3D scene understanding models. Additionally, we contribute a novel depth-aware scaling method, DeptS, that can effectively calibrate the uncertainty estimates of 3D scene understanding models.
3D Robustness. Several recent studies have been presented to benchmark the robustness of 3D perception models for driving applications. One line of research aims to probe the model’s 3D robustness against common corruptions from point clouds [55, 39, 38, 93, 12, 124, 115], depth [59, 32, 56], multi-view images [116, 46, 15], etc. Other related works explore the 3D robustness from aspects like sensor failures [126, 35] and adversarial attacks [130, 117]. In this work, different from all previous pursuits, we target another important direction on this relevant topic. As far as we know, Calib3D establishes the first 3D robustness benchmark from the uncertainty estimation perspective. Extensive experiments of 3D scene understanding models on diverse datasets are conducted, which could shed light on the future development of practical and reliable 3D scene understanding models.
3 Calib3D
A typical point cloud data $\mathcal{P}=\{\mathbf{p}_{i},q_{i}|i=0,1,...,N-1\}$ consists of a total of $N$ points acquired by the sensor, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ holds the Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$ and $q_{i}\in\mathbb{R}^{1}$ is the sensor reflection value, e.g., the laser intensity. For a learning-based system, there is an associated semantic label $\mathcal{Y}=\{y_{i}|i=0,1,...,N-1\}$ for each point in $\mathcal{P}$, where $y_{i}$ takes values from $S$ pre-defined semantic classes, depending on the specific label mapping. The random variables $\mathcal{P}$ and $\mathcal{Y}$ follow a ground truth joint distribution $\pi(\mathcal{P},\mathcal{Y})=\pi(\mathcal{Y}|\mathcal{P})\pi(\mathcal{P})$.
Problem Formulation. Let $h(\cdot)$ be a 3D semantic segmentation model that takes point cloud $\mathcal{P}$ as the input and outputs the class predictions $\hat{Y}=\{\hat{y}_{i}|i=0,1,...,N-1\}$ and the corresponding confidence scores $\hat{C}=\{\hat{c}_{i}|i=0,1,...,N-1\}$, i.e., $h(\mathcal{P})=(\hat{Y},\hat{C})$. Our goal is two-fold: (1) we would like to measure how well the 3D semantic segmentation model delivers the uncertainty estimates when making predictions, and (2) we anticipate obtaining a well-calibrated 3D semantic segmentation model that makes accurate predictions when the associated confidence scores are high. In theory, a perfect model calibration is defined as $\mathbb{P}(\hat{y}_{i}=y_{i}|\hat{c}_{i}=c)=c$, where $c\in[0,1]$ is the expected confidence value.
Objective. To better cater to real-world requirements, we resort to the non-probabilistic [^3] In the deep learning context, the non-probabilistic output $\mathbf{z}_{i}$ is often known as logits. output $\mathcal{Z}=\{\mathbf{z}_{i}|i=0,1,...,N-1\}$ from the 3D semantic segmentation model for calibration, without altering the model’s accuracy. The predicted probability $\hat{c}_{i}$ can be derived from $\mathbf{z}_{i}$ using a Softmax function, i.e., $\hat{c}_{i}=\sigma(\mathbf{z}_{i})$, with $\sigma(\cdot)$ denoting the Softmax operation. The overall objective is to produce a calibrated probability $\hat{v}_{i}$ for each point in $\mathcal{P}$, based on $\hat{y}_{i}$, $\hat{c}_{i}$, and $\mathbf{z}_{i}$.
3.1 Calibration Metrics
Expected Calibration Error (ECE). Guo et al. [36] proposed the ECE metric to assess the confidence calibration of a given neural network. Specifically, ECE measures the difference in expectation between confidence and accuracy:
$e_{ece}=\mathbb{E}_{\hat{c}_{i}}[~{}|~{}\mathbb{P}(\hat{y}_{i}=y_{i}~{}|~{}\hat{c}_{i}=c)-c~{}|~{}]~{}.$ (1)
Based on the definition, a perfectly calibrated model has an ECE value of zero.
ECE for 3D Scene Understanding. In practice, we approximate the value of Eq. 1 by quantizing continuous variables into equally interval probability bins. Different from the conventional image classification task [36, 71], we treat each of the $N$ points [^4] It is worth noting that point clouds often contain an unequal number of points, i.e., take different $N$ values, due to acquisitions; we omit such a difference for simplicity. in $\mathcal{P}$ as unique samples. Assuming a total of $\tilde{N}$ point clouds in a dataset, we first calculate the ECE score of each point cloud and then average across all point clouds. Such a statistical binning takes the weighted average of the accuracy/confidence difference of each bin in each sample as follows:
$\hat{e}_{ece}=\frac{1}{\tilde{N}}\sum^{\tilde{N}}_{\tilde{n}=1}\sum^{M}_{m=1}\frac{|B_{m}|}{N}\left|~{}\text{acc}(B_{m})-\text{conf}(B_{m})~{}\right|~{},$ (2)
where $M$ denotes the number of bins used for quantization. $B_{m}$ denotes the set of samples falling into the $m$-th bin. The difference between $\text{acc}(\cdot)$ and $\text{conf}(\cdot)$ is also known as the calibration gap and can be interpreted using reliability diagrams [22, 78, 36]. In the next section, we review the most popular post-hoc methods that have been widely used for calibration in the 2D community.
3.2 Calibration Methods
Temperature Scaling (TempS). As has been widely verified in theory and practice, a simple extension of the Platt scaling [86, 78] is effective in improving the model calibration. Following [36], we use a single temperature parameter $T>0$ for all $S$ classes to re-scale the non-probabilistic output $\mathbf{z}_{i}$, that is:
$\hat{v}_{i}^{\text{TempS}}=\max_{s}\sigma(\frac{\mathbf{z}_{i}}{T})^{(s)}~{},$ (3)
where $\sigma(\cdot)$ denotes the Softmax function. The $\max(\cdot)$ operation takes the maximum value at the $s$-th position over $S$ semantic classes. Conventionally, $T$ is learned by minimizing the log-likelihood loss on a validation set.
Logistic Scaling (LogiS). A more flexible version of the temperature scaling adopts more complex transformations during re-scaling. Guo et al. [36] proposed to use the logistic regression to adjust the non-probabilistic output $\mathbf{z}_{i}$, that is:
$\hat{v}_{i}^{\text{LogiS}}=\max_{s}\sigma(\mathbf{W}\cdot\mathbf{z}_{i}+\mathbf{b})^{(s)}~{},$ (4)
where $\mathbf{W}$ and $\mathbf{b}$ are optimized based on the negative log-likelihood loss on a validation set. In this work, we adopt the vector scaling variant of logistic scaling, where the parameter $\mathbf{W}$ is restricted to a diagonal matrix.
Dirichlet Scaling (DiriS). Assuming the model’s outputs follow a Dirichlet distribution (rather than just single probability values), Kull et al. [61] further derived the Dirichlet scaling from logistic scaling, that is:
$\hat{v}_{i}^{\text{DiriS}}=\max_{s}\sigma\left(\mathbf{W}\cdot\log(\sigma(\mathbf{z}_{i}))+\mathbf{b}\right)^{(s)}~{},$ (5)
Figure 2: Depth-correlated patterns in a ±$\pm 50$m LiDAR scene from SemanticKITTI [7]. (a) Ground truth. (b) Point-wise ECE scores. (c) Point-wise entropy scores.
Figure 3: Depth-wise confidence and accuracy statistics of uncalibrated (UnCal), temperature scaling (TempS), meta-calibration (MetaC), and our proposed depth-aware scaling (DeptS) methods.
where $\mathbf{W}$ and $\mathbf{b}$ are parameters for a linear parameterization of the predicted probability $\sigma(\mathbf{z}_{i})$, and, similar to Eq. 3 and Eq. 4, $\mathbf{W}$ and $\mathbf{b}$ can be optimized based on the negative log-likelihood loss on a validation set.
Meta-Calibration (MetaC). Ma et al. [71] proposed to combine a base calibrator (e.g., temperature scaling) with a bipartite ranking model for better calibration performance. Specifically, the prediction entropy calculated via $\hat{c}_{i}$ is used to select calibrators; the base calibrator will be used if the entropy is lower than a threshold, and, on the contrary, the predicted output will take random values. Theoretical analyses on high-probability bounds w.r.t. mis-coverage rate and coverage accuracy are presented in [71]. In practice, this is formulated as:
$\hat{v}_{i}^{\text{MetaC}}=\left\{\begin{aligned} S^{-1}~{},&~{}~{}~{}\text{if}-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{T})^{(s)}~{},&~{}~{}~{}\text{otherwise}\end{aligned}~{},\right.$ (6)
where temperature $T>0$ is learned by minimizing the log-likelihood loss on a validation set. $\eta$ is a hand-crafted threshold for filtering high-entropy predictions. It is worth noting that meta-calibration, albeit proven effective in previous literature, will inevitably lose accuracy preservation. The first condition in Eq. 6 introduces randomness to model predictions, which is likely to be impractical regarding real-world, safety-critical applications, e.g., 3D scene understanding.
3.3 DeptS: Depth-Aware Scaling for Better 3D Calibration
Observations. While previous calibration methods [86, 36, 61, 71] have shown appealing calibration performance on image-based perception tasks, their effectiveness on 3D data remains unknown. Different from RGB images, the point cloud data are unordered and texture-less, which inherits extra difficulties in feature learning [31, 29, 6]. As shown in Fig. 2, we observe a close correlation among calibration error, prediction entropy, and depth – an inherent 3D information that can be calculated based on the Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$.
Depth Correlations. To further consolidate this finding, we conduct a quantitative analysis of the relation between the calibration error and depth, which is depicted in Fig. 3. We calculate the statistics of confidence and accuracy scores of all LiDAR points and then split them into 10 bins based on their depth values, where each bin corresponds to a 5-meter range. We notice from the uncalibrated result that LiDAR points with large depth values (i.e., at the middle-to-far regions of the ego-vehicle) tend to have low accuracy. However, the confidence scores of the uncalibrated model do not decrease correspondingly, which in turn causes higher calibration errors. This motivates us to design a calibration method that can resolve the over-confidence issue for LiDAR points with large depths.
Depth-Aware Scaling. To fulfill the above pursuit, we propose a straightforward yet effective depth-aware scaling (DeptS) method for better calibrating 3D scene understanding models. DeptS consists of two base calibrators which are selectively used based on the prediction entropy calculated via $\hat{c}_{i}$, that is:
$\hat{v}_{i}^{\text{DeptS}}=\left\{\begin{aligned} \max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{1}})^{(s)}~{},&~{}~{}~{}\text{if}-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{\alpha\cdot T_{2}})^{(s)}~{},&~{}~{}~{}\text{otherwise}\end{aligned}~{},\right.$ (7)
where $T_{1}$ and $T_{2}$ are the temperature parameters and satisfy $T_{1}>T_{2}$. $\eta$ is a threshold for filtering high-entropy predictions. A larger entropy value means a higher probability of being misclassified, and misclassified samples tend to encounter the over-confidence problem [13]. Therefore, we use a large $T_{1}$ to smooth the logits distribution, which will in turn reduce the confidence score.
To address the issue of over-confidence for LiDAR points with large depths, we set a depth-correlation coefficient $\alpha$ and use it to re-weight the temperature parameters. The overall process is formulated as follows:
$\alpha=k_{1}\cdot d_{i}+k_{2}~{},~{}~{}~{}~{}d_{i}=\sqrt{(p^{x}_{i})^{2}+(p^{y}_{i})^{2}+(p^{z}_{i})^{2}}~{},$ (8)
where $k_{1}$ and $k_{2}$ are two learnable parameters and satisfy $k_{1}>0$. $d_{i}$ is the depth that is calculated based on the Cartesian coordinates. In such cases, the LiDAR points with large depth values will have large $\alpha$ values, which are then used to reduce the corresponding confidence score. This in turn mitigates the over-confidence issue for points in the middle-to-far regions. The comparison between our method and the uncalibrated model in Fig. 3 shows the effectiveness of our depth-aware confidence adjustment design. As we will discuss more concretely in the following sections, DeptS contributes a stable improvement in calibrating 3D scene understanding models across a diverse spectrum of 3D datasets.
3.4 Calib3D Configurations
Our study serves as an early attempt at understanding the predictive preferences of 3D scene understanding models. To consolidate our findings and observations on this topic, we make efforts from the following two aspects in Calib3D.
Aleatoric Uncertainty. 3D data are known to be highly heterogeneous due to different sensor acquisitions, placements, scene conditions, annotation protocols, pre-processing, etc. A learning-based system trained on different data sources tends to exhibit diverse predictive confidence and accuracy, especially under measurement noises. To probe such aleatoric uncertainties, a total of 10 popular 3D datasets are used in Calib3D, including 1nuScenes [29], 2SemanticKITTI [7], 3Waymo Open [99], 4SemanticPOSS [83], 5Synth4D [95], 6SemanticSTF [115], 7ScribbleKITTI [107], 8S3DIS [3], and the 9nuScenes-C and 10SemanticKITTI-C datasets from the Robo3D benchmark [55]. We hope such a comprehensive benchmark study could lay a solid foundation for the future development of reliable 3D scene understanding and beyond. Due to space limits, kindly refer to the Appendix for additional configuration details of these datasets.
Epistemic Uncertainty. The diverse spectrum of 3D scene understanding models introduces diverse factors that cope with the model uncertainty. Tailored to understand the model architecture and configuration details, Calib3D encompasses a total of 28 state-of-the-art models that have achieved promising accuracy on standard benchmarks. Based on the use of LiDAR representations, we categorize these models into five groups, i.e., 1range view [75, 21, 131, 16, 2, 54, 122], 2bird’s eye view (BEV) [129], 3voxel [18, 133, 20], 4multi-view fusion [119, 123, 101, 64, 91, 67], and 5point-based [103, 128, 112, 89, 90, 111, 121] models. Additionally, we also investigate the impact of recent 3D data augmentation techniques [113, 58, 122] and different sparse convolution backends [18, 101, 20], exhibiting design choices that matter for delivering accurate uncertainty estimates. Due to space limits, kindly refer to the Appendix for additional details.
4 Experiments
4.1 Settings
Implementation Details. We establish the Calib3D benchmark using the popular MMDetection3D [19] and OpenPCSeg [66] codebase. A total of 28 models and 10 datasets are covered in this study. We follow the default configurations when training the models, including the optimizer, learning rate, scheduler, number of training epochs, etc. The common 3D data augmentations used are random rotation, flipping, scaling, and jittering. As for the implementation of calibration methods, we follow the conventional setups from prior works. We calculate the statistics of the predictive entropy of correct/incorrect predictions and select the boundary value as the entropy threshold. We adopt a unified configuration when training the proposed DeptS and previous post-hoc calibration methods [36, 61, 110, 71]. All methods are trained for 20 epochs with a batch size of 8 and the AdamW optimizer [69]. The learning rate is set to $1$e33-3- 3 and the weight decay is $1$e66-6- 6. We use four GPUs for both training and evaluation.
Table 1: The expected calibration error (ECE, the lower the better) of state-of-the-art 3D scene understanding models on nuScenes [29] and SemanticKITTI [7]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our proposed depth-aware scaling calibrations, respectively.
Method Modal nuScenes [29] SemanticKITTI [7]
UnCal TempS LogiS DiriS MetaC DeptS UnCal TempS LogiS DiriS MetaC DeptS
RangeNet++ [75] Range $\bullet$ $4.57\%$ $2.74\%$ $2.79\%$ $2.73\%$ $2.78\%$ $2.61\%$ $4.01\%$ $3.12\%$ $3.16\%$ $3.59\%$ $2.38\%$ $2.33\%$
SalsaNext [21] Range $\bullet$ $3.27\%$ $2.59\%$ $2.58\%$ $2.57\%$ $2.52\%$ $2.42\%$ $5.37\%$ $4.29\%$ $4.31\%$ $4.11\%$ $3.35\%$ $3.19\%$
FIDNet [131] Range $\bullet$ $4.89\%$ $3.35\%$ $2.89\%$ $2.61\%$ $4.55\%$ $4.33\%$ $5.89\%$ $4.04\%$ $4.15\%$ $3.82\%$ $3.25\%$ $3.14\%$
CENet [16] Range $\bullet$ $4.44\%$ $2.47\%$ $2.53\%$ $2.58\%$ $2.70\%$ $2.44\%$ $5.95\%$ $3.93\%$ $3.79\%$ $4.28\%$ $3.31\%$ $3.09\%$
RangeViT [2] Range $\bullet$ $2.52\%$ $2.50\%$ $2.57\%$ $2.56\%$ $2.46\%$ $2.38\%$ $5.47\%$ $3.16\%$ $4.84\%$ $8.80\%$ $3.14\%$ $3.07\%$
RangeFormer [54] Range $\bullet$ $2.44\%$ $2.40\%$ $2.41\%$ $2.44\%$ $2.27\%$ $2.15\%$ $3.99\%$ $3.67\%$ $3.70\%$ $3.69\%$ $3.55\%$ $3.30\%$
FRNet [122] Range $\bullet$ $2.27\%$ $2.24\%$ $2.22\%$ $2.28\%$ $2.22\%$ $2.17\%$ $3.46\%$ $3.53\%$ $3.54\%$ $3.49\%$ $2.83\%$ $2.75\%$
PolarNet [129] BEV $\bullet$ $4.21\%$ $2.47\%$ $2.54\%$ $2.59\%$ $2.56\%$ $2.45\%$ $2.78\%$ $3.54\%$ $3.71\%$ $3.70\%$ $2.67\%$ $2.59\%$
MinkUNet18 [18] Voxel $\bullet$ $2.45\%$ $2.34\%$ $2.34\%$ $2.42\%$ $2.29\%$ $2.23\%$ $3.04\%$ $3.01\%$ $3.08\%$ $3.30\%$ $2.69\%$ $2.63\%$
MinkUNet34 [18] Voxel $\bullet$ $2.50\%$ $2.38\%$ $2.38\%$ $2.53\%$ $2.32\%$ $2.24\%$ $4.11\%$ $3.59\%$ $3.62\%$ $3.63\%$ $2.81\%$ $2.73\%$
Cylinder3D [133] Voxel $\bullet$ $3.19\%$ $2.58\%$ $2.62\%$ $2.58\%$ $2.39\%$ $2.29\%$ $5.49\%$ $4.36\%$ $4.48\%$ $4.42\%$ $3.40\%$ $3.09\%$
SpUNet18 [20] Voxel $\bullet$ $2.58\%$ $2.41\%$ $2.46\%$ $2.59\%$ $2.36\%$ $2.25\%$ $3.77\%$ $3.47\%$ $3.44\%$ $3.61\%$ $3.37\%$ $3.21\%$
SpUNet34 [20] Voxel $\bullet$ $2.60\%$ $2.52\%$ $2.47\%$ $2.66\%$ $2.41\%$ $2.29\%$ $4.41\%$ $4.33\%$ $4.34\%$ $4.39\%$ $4.20\%$ $4.11\%$
RPVNet [119] Fusion $\bullet$ $2.81\%$ $2.70\%$ $2.73\%$ $2.79\%$ $2.68\%$ $2.60\%$ $4.67\%$ $4.12\%$ $4.23\%$ $4.26\%$ $4.02\%$ $3.75\%$
2DPASS [123] Fusion $\bullet$ $2.74\%$ $2.53\%$ $2.51\%$ $2.51\%$ $2.62\%$ $2.46\%$ $2.32\%$ $2.35\%$ $2.45\%$ $2.30\%$ $2.73\%$ $2.27\%$
SPVCNN18 [101] Fusion $\bullet$ $2.57\%$ $2.44\%$ $2.49\%$ $2.54\%$ $2.40\%$ $2.31\%$ $3.46\%$ $2.90\%$ $3.07\%$ $3.41\%$ $2.36\%$ $2.32\%$
SPVCNN34 [101] Fusion $\bullet$ $2.61\%$ $2.49\%$ $2.54\%$ $2.61\%$ $2.37\%$ $2.28\%$ $3.61\%$ $3.03\%$ $3.07\%$ $3.10\%$ $2.99\%$ $2.86\%$
CPGNet [64] Fusion $\bullet$ $3.33\%$ $3.11\%$ $3.17\%$ $3.15\%$ $3.07\%$ $2.98\%$ $3.93\%$ $3.81\%$ $3.83\%$ $3.78\%$ $3.70\%$ $3.59\%$
GFNet [91] Fusion $\bullet$ $2.88\%$ $2.71\%$ $2.70\%$ $2.73\%$ $2.55\%$ $2.41\%$ $3.07\%$ $3.01\%$ $2.99\%$ $3.05\%$ $2.88\%$ $2.73\%$
UniSeg [67] Fusion $\bullet$ $2.76\%$ $2.61\%$ $2.63\%$ $2.65\%$ $2.45\%$ $2.37\%$ $3.93\%$ $3.73\%$ $3.78\%$ $3.67\%$ $3.51\%$ $3.43\%$
KPConv [103] Point $\bullet$ $3.37\%$ $3.27\%$ $3.34\%$ $3.32\%$ $3.28\%$ $3.20\%$ $4.97\%$ $4.88\%$ $4.90\%$ $4.91\%$ $4.78\%$ $4.68\%$
PIDS1.25× [128] Point $\bullet$ $3.46\%$ $3.40\%$ $3.43\%$ $3.41\%$ $3.37\%$ $3.28\%$ $4.77\%$ $4.65\%$ $4.66\%$ $4.64\%$ $4.57\%$ $4.49\%$
PIDS2.0× [128] Point $\bullet$ $3.53\%$ $3.47\%$ $3.49\%$ $3.51\%$ $3.34\%$ $3.27\%$ $4.91\%$ $4.83\%$ $4.72\%$ $4.89\%$ $4.66\%$ $4.47\%$
PTv2 [112] Point $\bullet$ $2.42\%$ $2.34\%$ $2.46\%$ $2.55\%$ $2.48\%$ $2.19\%$ $4.95\%$ $4.78\%$ $4.71\%$ $4.94\%$ $4.69\%$ $4.62\%$
WaffleIron [89] Point $\bullet$ $4.01\%$ $2.65\%$ $3.06\%$ $2.59\%$ $2.54\%$ $2.46\%$ $3.91\%$ $2.57\%$ $2.86\%$ $2.67\%$ $2.58\%$ $2.51\%$
Table 2: The expected calibration error (ECE) and segmentation accuracy (mIoU) of state-of-the-art 3D scene understanding models on six heterogeneous benchmarks. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
Dataset Type Method Modal UnCal TempS LogiS DiriS MetaC DeptS mIoU
Waymo Open [99] High-Res PolarNet [129] BEV $\bullet$ $3.92\%$ $1.93\%$ $1.90\%$ $1.91\%$ $2.39\%$ $1.84\%$ $58.33\%$
MinkUNet [18] Voxel $\bullet$ $1.70\%$ $1.70\%$ $1.74\%$ $1.76\%$ $1.69\%$ $1.59\%$ $68.67\%$
SPVCNN [101] Fusion $\bullet$ $1.81\%$ $1.79\%$ $1.80\%$ $1.88\%$ $1.74\%$ $1.69\%$ $68.86\%$
SemanticPOSS [83] Dynamic PolarNet [129] BEV $\bullet$ $4.24\%$ $8.09\%$ $7.81\%$ $8.30\%$ $5.35\%$ $4.11\%$ $52.11\%$
MinkUNet [18] Voxel $\bullet$ $7.22\%$ $7.44\%$ $7.36\%$ $7.62\%$ $5.66\%$ $5.48\%$ $56.32\%$
SPVCNN [101] Fusion $\bullet$ $8.80\%$ $6.53\%$ $6.91\%$ $7.41\%$ $4.61\%$ $3.98\%$ $53.51\%$
SemanticSTF [115] Weather PolarNet [129] BEV $\bullet$ $5.76\%$ $4.94\%$ $4.49\%$ $4.53\%$ $4.17\%$ $4.12\%$ $51.26\%$
MinkUNet [18] Voxel $\bullet$ $5.29\%$ $5.21\%$ $4.96\%$ $5.10\%$ $4.78\%$ $4.72\%$ $50.22\%$
SPVCNN [101] Fusion $\bullet$ $5.85\%$ $5.53\%$ $5.16\%$ $5.05\%$ $5.12\%$ $4.97\%$ $51.73\%$
ScribbleKITTI [107] Scribble PolarNet [129] BEV $\bullet$ $4.65\%$ $4.59\%$ $4.56\%$ $4.55\%$ $3.25\%$ $3.09\%$ $55.22\%$
MinkUNet [18] Voxel $\bullet$ $7.97\%$ $7.13\%$ $7.29\%$ $7.21\%$ $5.93\%$ $5.74\%$ $59.87\%$
SPVCNN [101] Fusion $\bullet$ $7.04\%$ $6.63\%$ $6.93\%$ $6.66\%$ $5.34\%$ $5.13\%$ $60.22\%$
Synth4D [95] Synthetic PolarNet [129] BEV $\bullet$ $1.68\%$ $0.93\%$ $0.75\%$ $0.72\%$ $1.54\%$ $0.69\%$ $85.63\%$
MinkUNet [18] Voxel $\bullet$ $2.43\%$ $2.72\%$ $2.43\%$ $2.05\%$ $4.01\%$ $2.39\%$ $69.11\%$
SPVCNN [101] Fusion $\bullet$ $2.21\%$ $2.35\%$ $1.86\%$ $1.70\%$ $3.44\%$ $1.67\%$ $69.68\%$
S3DIS [3] Indoor PointNet++ [90] Point $\bullet$ $9.13\%$ $8.36\%$ $7.83\%$ $8.20\%$ $6.93\%$ $6.79\%$ $56.96\%$
DGCNN [111] Point $\bullet$ $6.00\%$ $6.23\%$ $6.35\%$ $7.12\%$ $5.47\%$ $5.39\%$ $54.50\%$
PAConv [121] Point $\bullet$ $8.38\%$ $5.87\%$ $6.03\%$ $5.98\%$ $4.67\%$ $4.57\%$ $66.60\%$
Figure 4: The reliability diagrams of visualized calibration gaps from CENet [16] on the val set of SemanticKITTI [7]. UnCal, TempS, MetaC, and DeptS denote the uncalibrated, temperature, meta, and our depth-aware scaling calibrations, respectively.
Benchmark Protocols. To ensure fair comparisons, we unify the model training and evaluation configuration during benchmarking. The models are trained on the official training split of each dataset and are evaluated on data from the official val split. We reproduce the originally reported performance of the benchmarked models whenever possible, without using any extra tricks including test time augmentation, model ensembling, or fine-tuning on validation data.
Evaluation Metrics. We use the expected calibration error (ECE) metric as depicted in Eq. 2 as the main indicator in our benchmark. Following conventions, we adopt the class-wise Intersection-over-Union (IoU) and the mean Intersection-over-Union (mIoU) to measure the 3D segmentation accuracy. For robustness probing, we adopt the corruption-wise IoU scores and the mean Resilience Rate (mRR) defined in Robo3D [55] to measure the 3D robustness. Kindly refer to the Appendix for more details on these metrics.
4.2 In-Domain Uncertainty
Automotive 3D Scenes. Tab. 1 displays the calibration errors of state-of-the-art 3D scene understanding models on nuScenes [29] and SemanticKITTI [7]. We find that these models are prone to be poorly calibrated, which undermines their applicability in safety-critical contexts. We also observe similar patterns on Waymo Open [99] and SemanticPOSS [83] in Tab. 2. The calibrated uncertainties from different calibration methods [36, 61, 110, 71] exhibit promising rescues to remedy such problems. As can be seen, a simple temperature scaling [36] is useful in improving calibration. Our DeptS sets up new states for all models across all datasets, which verifies the efficacy of the proposed depth-aware scaling. A similar conclusion can be derived from Fig. 1, where DeptS holistically improves the uncertainty estimation in different regions from the LiDAR scenes.
Adverse Weather Conditions. The results of PolarNet [129], MinkUNet [18], and SPVCNN [101] on the SemanticSTF [115] dataset in Tab. 2 further showcase the importance of conducting network calibration in 3D scene understanding. Models under weather conditions fail to deliver accurate uncertainty estimates, posing challenges for real-world usage. Effective calibration using DeptS tends to provide a more reliable uncertainty estimation for safety-critical applications.
Synthetic LiDAR Data. The results on the Synth4D [95] dataset in Tab. 2 reveal the fact that models trained on synthetic data are likely to have lower calibration errors. This is because the point clouds generated by simulation engines are less complex than real-world cases. Extra attention should be drawn when adapting models trained using synthetic data to real-world applications.
Sparse Annotations. From Tab. 2, we find that models trained using weak supervisions, such as the line scribbles in ScribbleKITTI [107], tend to yield much higher calibration errors. Different from training using dense annotations, models can only learn limited knowledge from weak supervision, which inevitably leads to higher predictive uncertainties. It is thus suggested to adopt calibration methods under such cases, where the calibration errors can be effectively reduced.
Indoor 3D Scenes. The last three rows of Tab. 2 show the calibration errors on PointNet++ [90], DGCNN [111], and PAConv [121] on S3DIS [3]. As can be seen, the indoor point clouds also struggle with aleatoric and epistemic uncertainties. It becomes evident that conducting network calibration is a crucial step in achieving high levels of robustness for 3D scene understanding. Similar to other use cases in Tab. 1 and Tab. 2, our proposed DeptS serves as an effective way to reduce the gaps between confidence and predictive accuracy.
Table 3: The expected calibration error (ECE) of MinkUNet [18] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
Type nuScenes-C SemanticKITTI-C
UnCal TempS LogiS DiriS MetaC DeptS UnCal TempS LogiS DiriS MetaC DeptS
Clean $\bullet$ $2.45\%$ $2.34\%$ $2.34\%$ $2.42\%$ $2.29\%$ $2.23\%$ $3.04\%$ $3.01\%$ $3.08\%$ $3.30\%$ $2.69\%$ $2.63\%$
Fog $\circ$ $5.52\%$ $5.42\%$ $5.49\%$ $5.43\%$ $4.77\%$ $4.72\%$ $12.66\%$ $12.55\%$ $12.67\%$ $12.48\%$ $11.08\%$ $10.94\%$
Wet Ground $\circ$ $2.63\%$ $2.54\%$ $2.54\%$ $2.64\%$ $2.55\%$ $2.52\%$ $3.55\%$ $3.46\%$ $3.54\%$ $3.72\%$ $3.33\%$ $3.28\%$
Snow $\circ$ $13.79\%$ $13.32\%$ $13.53\%$ $13.59\%$ $11.37\%$ $11.31\%$ $7.10\%$ $6.96\%$ $6.95\%$ $7.26\%$ $5.99\%$ $5.63\%$
Motion Blur $\circ$ $9.54\%$ $9.29\%$ $9.37\%$ $9.01\%$ $8.32\%$ $8.29\%$ $11.31\%$ $11.16\%$ $11.24\%$ $12.13\%$ $9.00\%$ $8.97\%$
Beam Missing $\circ$ $2.58\%$ $2.48\%$ $2.49\%$ $2.57\%$ $2.53\%$ $2.47\%$ $2.87\%$ $2.83\%$ $2.84\%$ $2.98\%$ $2.83\%$ $2.79\%$
Crosstalk $\circ$ $13.64\%$ $13.00\%$ $12.97\%$ $13.44\%$ $9.98\%$ $9.73\%$ $4.93\%$ $4.83\%$ $4.86\%$ $4.81\%$ $3.54\%$ $3.48\%$
Incomplete Echo $\circ$ $2.44\%$ $2.33\%$ $2.33\%$ $2.42\%$ $2.32\%$ $2.21\%$ $3.21\%$ $3.19\%$ $3.25\%$ $3.48\%$ $2.84\%$ $2.19\%$
Cross Sensor $\circ$ $4.25\%$ $4.15\%$ $4.20\%$ $4.28\%$ $4.06\%$ $3.20\%$ $3.15\%$ $3.13\%$ $3.18\%$ $3.43\%$ $3.17\%$ $2.96\%$
Average $\bullet$ $6.78\%$ $6.57\%$ $6.62\%$ $6.67\%$ $5.74\%$ $5.56\%$ $6.10\%$ $6.01\%$ $6.07\%$ $6.29\%$ $5.22\%$ $5.03\%$
Reliability Diagrams. As mentioned in Sec. 3.1, the calibration gaps can be well displayed using reliability diagrams. Fig. 4 demonstrates again the superiority of DeptS; our method effectively reduces the calibration gaps (as depicted in red areas), delivering more accurate uncertainty estimates in practice than prior calibration methods [36, 71]. Additional reliability diagrams are in the Appendix.
Table 4: Ablation study on the uncertainty of 3D segmentation networks with different capacities (# of parameters).
MinkUNet UnCal TempS MetaC DeptS mIoU
14×14$\times$14 ×Layer $\bullet$ $2.25\%$ $2.21\%$ $2.19\%$ $2.08\%$ $73.48\%$
18×18$\times$18 ×Layer $\bullet$ $2.45\%$ $2.34\%$ $2.29\%$ $2.23\%$ $76.19\%$
34×34$\times$34 ×Layer $\bullet$ $2.50\%$ $2.38\%$ $2.32\%$ $2.22\%$ $76.99\%$
50×50$\times$50 ×Layer $\bullet$ $2.56\%$ $2.41\%$ $2.39\%$ $2.30\%$ $77.70\%$
101×101$\times$101 ×Layer $\bullet$ $2.60\%$ $2.46\%$ $2.35\%$ $2.20\%$ $79.69\%$
Table 5: Ablation study on the uncertainty of 3D segmentation networks with different 3D data augmentation methods.
Augment UnCal TempS MetaC DeptS mIoU
Common $\bullet$ $2.45\%$ $2.34\%$ $2.29\%$ $2.23\%$ $76.19\%$
PolarMix $\bullet$ $2.39\%$ $2.35\%$ $2.30\%$ $2.20\%$ $76.19\%$
LaserMix $\bullet$ $2.22\%$ $2.21\%$ $2.18\%$ $2.15\%$ $76.39\%$
FrustumMix $\bullet$ $2.27\%$ $2.26\%$ $2.25\%$ $2.21\%$ $76.43\%$
Combo $\bullet$ $2.21\%$ $2.21\%$ $2.23\%$ $2.18\%$ $77.15\%$
Table 6: Ablation study on the uncertainty of CENet [16] with different # of range view cells on SemanticKITTI [7].
# of Cells UnCal TempS MetaC DeptS mIoU
$64\times 512$bullet $5.65\%$ $4.01\%$ $3.16\%$ $3.09\%$ $60.92\%$
$64\times 1024$bullet $5.88\%$ $4.04\%$ $3.24\%$ $3.16\%$ $62.04\%$
$64\times 2048$bullet $5.95\%$ $3.93\%$ $3.21\%$ $3.10\%$ $61.18\%$
$64\times 3072$bullet $6.00\%$ $3.45\%$ $2.85\%$ $2.71\%$ $60.66\%$
$64\times 4096$bullet $6.21\%$ $3.19\%$ $2.90\%$ $2.73\%$ $58.68\%$
Table 7: Ablation study on the uncertainty of MinkUNet-18 [18] with different voxel sizes (cubic shape) on nuScenes [29].
Voxel Size UnCal TempS MetaC DeptS mIoU
$0.05$ meter3 $\bullet$ $2.32\%$ $2.30\%$ $2.28\%$ $2.23\%$ $71.59\%$
$0.07$ meter3 $\bullet$ $2.34\%$ $2.28\%$ $2.27\%$ $2.21\%$ $75.14\%$
$0.10$ meter3 $\bullet$ $2.45\%$ $2.34\%$ $2.29\%$ $2.23\%$ $76.19\%$
$0.15$ meter3 $\bullet$ $2.48\%$ $2.43\%$ $2.28\%$ $2.21\%$ $75.92\%$
$0.20$ meter3 $\bullet$ $2.68\%$ $2.60\%$ $2.36\%$ $2.25\%$ $75.53\%$
4.3 Domain-Shift Uncertainty
In addition to the in-domain scenarios, we also investigate the uncertainty estimates under the more challenging domain-shift conditions. Following the out-of-domain (OoD) configurations in Robo3D [55], the 3D scene understanding models are trained using in-domain data and are tested under OoD conditions.
Common Corruptions. The 3D data acquired from real-world scenes tend to incorporate inherent measurement noises and variants. The first four rows in Tab. 3 display such issues; the models suffer from much higher calibration errors under corruptions caused by adverse weather conditions, including fog, wet ground, and snow. The degradation from motion blur further highlights the necessity of conducting network calibration for reliable 3D scene understanding.
Sensor Failures. The beam missing, crosstalk, incomplete echo, and cross sensor scenarios in Tab. 3 reveal the vulnerability of existing 3D scene understanding models to diverse sensor failures. Compared to prior calibration methods [36, 61, 110, 71], DeptS is more stable in delivering reliable uncertainty estimates under these OoD conditions. We believe this feature is crucial towards achieving a more reliable 3D scene understanding, especially for safety-critical applications.
4.4 Ablation Study
In this section, we study several important configurations that coped closely with current 3D scene understanding research. To control variables, unless otherwise specified, we use a MinkUNet-18 [18] model with voxel size of $0.10$ m3, common augmentations, and TorchSparse [100, 102] backend on the nuScenes [29] dataset, throughout this ablation study.
Network Capacity. As has been proven in prior studies [36, 110, 71], large 2D models tend to be less calibrated than small ones. From Sec. 4.2, we observed a similar pattern for 3D scene understanding models. We find that models with fewer parameters yield lower calibration errors, albeit being less accurate. Such a phenomenon poses concerns about the development of large 3D models for safety-critical applications. Special attention should be drawn when designing models with larger capacities since they are prone to be less calibrated in practice.
3D Data Augmentations. Recent advancements in 3D data augmentations have exhibited superior 3D segmentation accuracy. In Sec. 4.2, we benchmark popular techniques, namely LaserMix [58], PolarMix [113], and FrustumMix [122], on their efficacy in uncertainty estimation. We observe large improvements in them in delivering reliable uncertain estimates, compared to their baselines.
3D Rasterization. Different rasterization resolutions tend to affect the model’s accuracy. In Sec. 4.2 and Sec. 4.2, we investigate the behaviors of models trained with different numbers of range view cells and different voxel sizes. We observe that an optimal segmentation accuracy tends to appear at a resolution that is not too high or too low. However, the calibration error poses a clear correlation with the 3D rasterization, where more range view cells or smaller voxel sizes lead to increased calibration errors and vice versa. Based on observation, it is thus suggested to take special care in configuring the training and evaluation resolutions of 3D scene understanding models with different LiDAR representations.
(a) ECE vs. mIoU
(b) SparseConv Backend
(c) LiDAR Modality
Figure 5: Ablation studies on (a) relationships between calibration error and intersection-over-union scores, (b) calibration errors of MinkUNet [18] using different sparse convolution backends, and (c) average calibration errors of different LiDAR representations.
Segmentation Accuracy. We find a distinct correlation between the calibration errors and the 3D segmentation accuracy, i.e., the mIoU scores, as shown in Fig. 5(a). Similar to the observation drawn in [109], we find that a model with higher task accuracy is likely to have a relatively lower calibration error.
SparseConv Backends. We compare the behaviors of MinkUNet [18] trained using different sparse convolution backends, namely MinkowskiEngine [18], SpConv [125, 20], and TorchSparse [100, 102], and display the results in Fig. 5(b). In a general sense, SpConv [125, 20] tends to yield a higher calibration error than the other two backends. Our DeptS shows better performance across three scenarios.
3D Representations. In Fig. 5(c), we calculate the average calibration errors from all models benchmarked in Tab. 1 and split them into groups based on the use of 3D representations. As can be seen, models with point and range view representations are less calibrated than other modalities. Fusion-based models exhibit superiority in general, which showcases their efficacy in real-world cases.
5 Conclusion
In this study, we have presented Calib3D, a pioneering benchmark aimed at evaluating the reliability and accuracy of uncertainty estimates of 3D scene understanding models. Through our comprehensive assessment of state-of-the-art models across diverse 3D datasets, we have illuminated the critical challenges these models face in delivering confident and accurate predictions, especially within the context of safety-critical applications. Our findings reveal a significant gap in the calibration capabilities of current 3D scene understanding models, which often exhibit commendable accuracy but fall short in aligning their confidence levels with their predictive accuracy. Addressing this crucial gap, we introduced DeptS, a depth-aware scaling method specifically tailored to enhance the calibration of 3D models. By dynamically adjusting the logits distribution based on depth-correlated temperature scaling, the calibration performance is significantly improved. We hope that Calib3D and DeptS will inspire further research and innovation in the field of reliable 3D scene understanding.
Appendix
6 Calib3D Benchmark
In this section, we elaborate on additional details about the proposed Calib3D benchmark, including basic configurations regarding the datasets (Sec. 6.1), models (Sec. 6.2), evaluation protocols (Sec. 6.3), and license (Sec. 6.4).
6.1 3D Datasets
The Calib3D benchmark encompasses a total of 10 popular datasets in the area of 3D scene understanding, with a diverse spectrum of dataset configurations regarding data collections, label mappings, and annotation protocols. Tab. 8 provides an overview of the datasets used in our benchmark. The key features of each dataset are summarized as follows.
  • nuScenes [29] is one of the most popular driving datasets in autonomous vehicle research, featuring multimodal data from Boston and Singapore. It contains $1000$ scenes with $1.1$ billion annotated LiDAR points acquired by a Velodyne HDL32E LiDAR sensor. In this work, we use the lidarseg subset of the Panoptic-nuScenes dataset, which provides point-wise class and instance labels across 16 merged semantic categories. For more information: https://www.nuscenes.org/nuscenes.
  • SemanticKITTI [7] offers $22$ densely labeled LiDAR sequences of urban street scenes, making it one of the most prevailing benchmarks for LiDAR-based semantic scene understanding. The point clouds are acquired by a Velodyne HDL-64E LiDAR sensor and are annotated with a total of 19 semantic categories. For more information: http://semantic-kitti.org.
  • Waymo Open Dataset (WOD) [99] is a large-scale dataset for autonomous driving. The 3D semantic segmentation subset of WOD comprises $1150$ scenes, which are further split into $798$ training, $202$ validation, and $150$ testing scenes, corresponding to $23691$ training scans, $5976$ validation scans, and $2982$ testing scans, respectively. The LiDAR scans are annotated across $22$ semantic categories. For more information: https://waymo.com/open.
  • SemanticPOSS [83] is constructed with a special focus on dynamic scenes. It includes $2988$ scans from a 40-beam Hesai Pandora LiDAR sensor, offering insights into scene dynamics at Peking University’s campus. For more information: https://www.poss.pku.edu.cn/semanticposs.
  • SemanticSTF [115] is built on the STF dataset [8]. It features $2076$ scans under various weather conditions in the real world, serving as a testbed for assessing model robustness. The point clouds are acquired by a Velodyne HDL64 S3D LiDAR sensor under snowy, foggy, and rainy scenarios. For more information: https://github.com/xiaoaoran/SemanticSTF.
  • ScribbleKITTI [107] is an extension of the SemanticKITTI [7] dataset. It introduces weakly-supervised annotations through line scribbles, offering a cost-effective labeling approach for $19130$ LiDAR scans, which are under the same data splits and semantic annotations of citebehley2019semanticKITTI. For more information: https://github.com/ouenal/scribblekitti.
  • Synth4D [95] was collected utilizing CARLA simulations [26]. The Synth4D-nuScenes subset contains about $20000$ labeled point clouds for testing model performance in virtual urban and rural scenes, where the label mappings are aligned with that of the nuScenes [29] dataset. For more information: https://github.com/saltoricristiano/gipso-sfouda.
  • S3DIS [3] is a comprehensive collection of point clouds for indoor spaces. It encompasses detailed scans from six large-scale indoor areas that include over 215 million points and covers more than 6,000 square meters. Each point in the dataset is annotated with one of several semantic labels corresponding to different object categories like walls, floors, chairs, tables, etc. For more information: http://buildingparser.stanford.edu/dataset.html.
  • nuScenes-C [55] is part of the 3D robustness benchmarks in Robo3D [55] and is built based on the nuScenes [29] dataset. It focuses on the 3D model’s out-of-distribution robustness against eight types of common corruptions, offering a platform for testing under diverse adverse conditions. For more information: https://github.com/ldkong1205/Robo3D.
  • SemanticKITTI-C [55] shares the same common corruption types with nuScenes-C and is built based on the SemanticKITTI [7] dataset. For more information: https://github.com/ldkong1205/Robo3D.
Table 8: Summary of 3D datasets encompassed in the Calib3D benchmark, including 1nuScenes [29], 2SemanticKITTI [7], 3Waymo Open [99], 4SemanticPOSS [83], 5SemanticSTF [115], 6ScribbleKITTI [107], 7Synth4D [95], 8S3DIS [3], and 9nuScenes-C and 10SemanticKITTI-C from the Robo3D benchmark [55]. Each dataset sheds light on a specific data acquisition and annotation protocol, such as different LiDAR sensor setups, adverse weather conditions, weak annotations, synthetic data, indoor scenes, and out-of-domain corruptions. The images are adopted from the original papers.
nuScenes SemanticKITTI Waymo Open SemanticPOSS SemanticSTF
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ScribbleKITTI Synth4D S3DIS nuScenes-C SemanticKITTI-C
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
6.2 3D Models
The Calib3D benchmark encompasses a total of 28 state-of-the-art models in the area of 3D scene understanding, with a diverse spectrum of LiDAR representations, network architectures, and pre-/post-processing. Tab. 9 provides a summary of the models used, including their LiDAR modalities and key features.
Table 9: Summary of 3D models encompassed in the Calib3D benchmark. We categorize models into five distinct groups, based on their LiDAR representations, i.e., 1range view, 2bird’s eye view, 3sparse voxel, 4multi-view fusion, and 5raw point. Each model sheds light on a specific network structure and model configuration.
Model Modality Key Feature Ref
RangeNet++ $\bullet$ Range View The first range view LiDAR segmentation model with a FCN structure [75]
SalsaNext $\bullet$ Range View Uncertainty-aware range view segmentation with dilation modules [21]
FIDNet $\bullet$ Range View Fully interpolation encoding for better range view post processing [131]
CENet $\bullet$ Range View Concise and efficient range view learning with unified model structure [16]
RangeViT $\bullet$ Range View Replace ResNet backbone with ViT for enhancing range view learning [2]
RangeFormer $\bullet$ Range View Combine RangeAug, RangePost, and RangeFormer for better results [54]
FRNet $\bullet$ Range View Frustum-range fusion & interpolation for scalable LiDAR segmentation [122]
PolarNet $\bullet$ Bird’s Eye View Point cloud embedding using polar coordinates for real-time processing [129]
MinkUNet18 $\bullet$ Sparse Voxel Highly efficient sparse convolution operators with cubic voxel grids [18]
MinkUNet34 $\bullet$ Sparse Voxel Enhanced MinkUNet structure with a larger segmentation backbone [18]
Cylinder3D $\bullet$ Sparse Voxel Cylindrical voxel representation for balanced LiDAR points encoding [133]
SpUNet18 $\bullet$ Sparse Voxel MinkUNet structure with SpConv operators for efficient 3D learning [20]
SpUNet34 $\bullet$ Sparse Voxel Enhanced SpUNet structure with a larger segmentation backbone [20]
RPVNet $\bullet$ Multi-View Fusion Multi-view fusion of range, point, and voxel for modality interactions [119]
2DPASS $\bullet$ Multi-View Fusion Distillation from images to enhance point cloud feature learning [123]
SPVCNN18 $\bullet$ Multi-View Fusion Efficient sparse point-voxel convolutions & a lightweight architecture [101]
SPVCNN34 $\bullet$ Multi-View Fusion Enhanced SPVCNN structure with a larger segmentation backbone [101]
CPGNet $\bullet$ Multi-View Fusion Cascade point-grid fusion & transformation consistency regularization [64]
GFNet $\bullet$ Multi-View Fusion Complementary geometric flow fusion of range and bird’s eye views [91]
UniSeg $\bullet$ Multi-View Fusion Unified multi-view representation learning and cross-view distillation [67]
KPConv $\bullet$ Raw Point Input Deformable convolutions for adaptive kernel-based geometry learning [103]
PIDS1.25× $\bullet$ Raw Point Input Joint point interaction-dimension search with varying point densities [128]
PIDS2.0× $\bullet$ Raw Point Input Enhanced PIDS structure with a larger segmentation backbone [128]
PTv2 $\bullet$ Raw Point Input Grouped vector attention & partition-based pooling using Transformers [112]
WaffleIron $\bullet$ Raw Point Input Update point features by combining multi-MLPs and dense 2D CNNs [89]
PointNet++ $\bullet$ Raw Point Input The first hierarchical network to direct operate on point clouds [90]
DGCNN $\bullet$ Raw Point Input Use graph convolution to dynamically update graph in feature space [111]
PAConv $\bullet$ Raw Point Input Dynamic kernel assembling to adjust convolutions with point positions [121]
6.3 Benchmark Protocols
In this work, to ensure fairness in comparisons, we adopt the following protocols in model evaluations:
  • All 3D scene understanding models are trained on the official training set of each 3D dataset, and evaluated on data from the official validation set. There is no overlap between training and evaluation data.
  • To reflect the original behavior of each 3D scene understanding model, we directly use public checkpoints whenever applicable, or re-train the model using its default configuration. The acknowledgments of public checkpoints and implementations are included in Sec. 11.
  • We notice that some models (and their public checkpoints) are enhanced using extra “tricks” on the validation/testing sets, such as test time augmentation, model ensembling, etc. To ensure fairness, we re-train such models to reflect their “clean” performance.
6.4 License
The Calib3D benchmark is released under the CC BY-NC-SA 4.0 license [^5] https://creativecommons.org/licenses/by-nc-sa/4.0/legalcode.en.. For licenses regarding the codebase used in the Calib3D benchmark, kindly refer to Sec. 11.1. For licenses regarding the 3D datasets used in the Calib3D benchmark, kindly refer to Sec. 11.2. For licenses regarding the model implementations used in the Calib3D benchmark, kindly refer to Sec. 11.3.
7 Additional Implementation Detail
In this section, we provide additional implementation details to help reproduce the key results shown in this work.
7.1 3D Model Training
Our Calib3D benchmark is constructed based on the popular MMDetection3D [19] and OpenPCSeg [66] codebase, as well as several standalone implementations that have not been integrated into MMDetection3D [19] and/or OpenPCSeg [66]. Most 3D models adopt a unified training configuration, including the number of training epochs, optimizer, and learning rate scheduler. We apply common 3D data augmentations in Cartesian space, including random flipping, rotation, scaling, and jittering. The 3D models are trained using eight GPUs with a batch size of 2. The number of epochs are set as $80$ for nuScenes and $50$ for SemanticKITTI, Waymo Open, SemanticPOSS, SemanticSTF, ScribbleKITTI, and Synth4D. For S3DIS, we follow the default setups as MMDetection3D [19]. For additional details, please refer to the corresponding codebase.
7.2 3D Model Evaluation
We evaluate the 3D models by following the conventional evaluation setups. As mentioned in Sec. 6.3, we do not use any extra “tricks” on the validation/testing sets, such as test time augmentation, model ensembling, etc.
7.3 PyTorch-Style ECE Calculation
To facilitate reproduction, we provide a PyTorch-style code snippet for calculating the expected calibration error (ECE) on point clouds in LABEL:lst:ece3d.
7.4 PyTorch-Style Implementation of DeptS
To facilitate reproduction, we provide a PyTorch-style code snippet of the proposed depth-aware scaling (DeptS) method in LABEL:lst:depts.
1import torch
2import torch.nn.functional as F
3
4def calculate_ece(logits, labels, ignore_index, n_bins=10):
5 valid_index = labels != ignore_index
6 logits, labels = logits[valid_index], labels[valid_index]
7
8 bin_bound = torch.linspace(0, 1, n_bins + 1)
9 lowers, uppers = bin_bound[:-1], bin_bound[1:]
10
11 softmaxes = F.softmax(logits, dim=1)
12 confs, preds = torch.max(softmaxes, 1)
13 accs = preds.eq(labels)
14
15 ece = torch.zeros(1)
16 for l, u in zip(lowers, uppers):
17 in_bin = confs.gt(l.item()) * confs.le(u.item())
18 prop_in_bin = in_bin.float().mean()
19 if prop_in_bin.item() > 0:
20 acc_in_bin = accs[in_bin].float().mean()
21 avg_conf_in_bin = confs[in_bin].mean()
22 ece += torch.abs(avg_conf_in_bin - acc_in_bin) * prop_in_bin
23
24 return ece.item()
Listing 1: PyTorch-style code snippet for calculating ECE scores on point clouds.
1import numpy as np
2import torch
3import torch.nn as nn
4
5class Depth_Aware_Scaling(nn.Module):
6
7 def __init__(self, threshold):
8 super(Depth_Aware_Scaling, self).__init__()
9 self.T1 = nn.Parameter(torch.ones(1))
10 self.T2 = nn.Parameter(torch.ones(1) * 0.9)
11 self.k = nn.Parameter(torch.ones(1) * 0.1)
12 self.b = nn.Parameter(torch.zeros(1))
13 self.alpha = 0.05
14 self.threshold = threshold
15 self.softmax = nn.Softmax(dim=-1)
16
17 def forward(self, logits, gt, xyz):
18 if self.training:
19 ind = torch.argmax(logits, axis=1) == gt
20 logits_pos, gt_pos = logits[ind], gt[ind]
21 logits_neg, gt_neg = logits[~ind], gt[~ind]
22
23 depth = torch.norm(xyz, p=2, dim=1)
24 depth_pos, depth_neg = depth[ind], depth[~ind]
25
26 s = np.random.randint(int(logits_pos.shape[0] * 1 / 3)) + 1
27 logits = torch.cat((
28 logits_neg, logits_pos[s:int(logits_pos.shape[0] / 2) + s]
29 ), 0)
30 gt = torch.cat((
31 gt_neg, gt_pos[s:int(logits_pos.shape[0] / 2) + s]
32 ), 0)
33 depth = torch.cat((
34 depth_neg, depth_pos[s:int(depth_pos.shape[0] / 2) + s]
35 ), 0)
36
37 prob = self.softmax(logits)
38
39 score = torch.sum(-prob * torch.log(prob), dim=-1)
40 cond_ind = score < self.threshold
41
42 cal_logits_1, cal_gt_1 = logits[cond_ind], gt[cond_ind]
43 cal_logits_2, cal_gt_2 = logits[~cond_ind], gt[~cond_ind]
44
45 depth_coff = self.k * depth + self.b
46 T1 = self.T1 * depth_coff[cond_ind].unsqueeze(dim=-1)
47 T2 = self.T2 * depth_coff[~cond_ind].unsqueeze(dim=-1)
48
49 cal_logits_1 = cal_logits_1 / T1
50 cal_logits_2 = cal_logits_2 / T2
51
52 cal_logits = torch.cat((cal_logits_1, cal_logits_2), 0)
53 cal_gt = torch.cat((cal_gt_1, cal_gt_2), 0)
54
55 else:
56 prob = self.softmax(logits)
57
58 score = torch.sum(-prob * torch.log(prob), dim=-1)
59 cond_ind = score < self.threshold
60
61 scaled_logits, scaled_gt = logits[cond_ind], gt[cond_ind]
62 inference_logits, inference_gt = logits[~cond_ind], gt[~cond_ind]
63
64 depth = torch.norm(xyz, p=2, dim=1).to(logits.device)
65 depth_coff = self.k * depth + self.b
66
67 T1 = self.T1 * depth_coff[cond_ind].unsqueeze(dim=-1)
68 T2 = self.T2 * depth_coff[~cond_ind].unsqueeze(dim=-1)
69
70 scaled_logits = scaled_logits / T1
71 inference_logits = inference_logits / T2
72
73 cal_logits = torch.cat((scaled_logits, inference_logits), 0)
74 cal_gt = torch.cat((scaled_gt, inference_gt), 0)
75
76 return cal_logits, cal_gt
Listing 2: PyTorch-style code snippet of the proposed depth-aware scaling (DeptS).
8 Additional Quantitative Result
In this section, we supplement additional quantitative results to better support the findings and conclusions drawn in the main body of this paper.
8.1 Reliability Diagrams
We provide additional reliability diagrams in Fig. 6 for a more comprehensive validation of the effectiveness of our method. As can be seen, the 3D models without proper calibration (UnCal) tend to suffer from huge confidence-accuracy gaps. This inevitably leads to potential impediments to the safe operation of 3D scene understanding systems in the real world. Our compared calibration methods show effectiveness in mitigating such issues. Compared to the previous calibration methods, our DeptS exhibits superior performance across a wide spectrum of scenarios. This can be credited to the depth-aware scaling operation which encourages a more consistent prediction in depth-correlated areas.
8.2 Domain-Shift Uncertainty Estimation
Enhancing the uncertainty estimation capability of handling challenging scenarios is crucial for the practical usage of 3D scene understanding systems. We supplement the domain-shift uncertainty estimation results of FRNet [122] and SPVCNN [101] in Tab. 10 and Tab. 11, respectively. Similar to the observations drawn in the main body of this paper, we find that 3D models are vulnerable under adverse conditions. The expected calibration errors are extremely high under “fog”, “motion blur”, “crosstalk”, and “cross sensor” corruptions, which are commonly occurring scenarios in the real world. Compared to previous calibration methods like temperature scaling and meta-calibration, our DeptS shows a more stable performance across different domain-shift scenarios. We believe such an ability will become more and more important in the future development of 3D scene understanding systems.
Table 10: The expected calibration error (ECE) of FRNet [122] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
Type nuScenes-C SemanticKITTI-C
UnCal TempS LogiS DiriS MetaC DeptS UnCal TempS LogiS DiriS MetaC DeptS
Clean $\bullet$ $2.27\%$ $2.24\%$ $2.22\%$ $2.28\%$ $2.22\%$ $2.17\%$ $3.46\%$ $3.53\%$ $3.54\%$ $3.49\%$ $2.83\%$ $2.75\%$
Fog $\circ$ $3.07\%$ $3.06\%$ $3.07\%$ $3.03\%$ $3.06\%$ $2.98\%$ $13.48\%$ $13.57\%$ $13.66\%$ $13.47\%$ $12.68\%$ $12.42\%$
Wet Ground $\circ$ $2.46\%$ $2.44\%$ $2.43\%$ $2.50\%$ $2.56\%$ $2.41\%$ $4.01\%$ $4.09\%$ $4.11\%$ $3.96\%$ $3.32\%$ $3.28\%$
Snow $\circ$ $3.50\%$ $3.42\%$ $3.53\%$ $3.60\%$ $2.93\%$ $2.78\%$ $7.28\%$ $7.39\%$ $7.49\%$ $7.51\%$ $6.65\%$ $6.63\%$
Motion Blur $\circ$ $33.74\%$ $33.48\%$ $33.15\%$ $32.15\%$ $30.62\%$ $28.43\%$ $5.93\%$ $6.03\%$ $6.08\%$ $6.55\%$ $5.04\%$ $4.92\%$
Beam Missing $\circ$ $2.52\%$ $2.51\%$ $2.50\%$ $2.58\%$ $2.91\%$ $2.48\%$ $2.71\%$ $2.71\%$ $2.72\%$ $2.71\%$ $2.40\%$ $2.36\%$
Crosstalk $\circ$ $2.40\%$ $2.39\%$ $2.36\%$ $2.38\%$ $2.72\%$ $2.35\%$ $20.87\%$ $21.16\%$ $21.03\%$ $19.84\%$ $15.36\%$ $14.79\%$
Incomplete Echo $\circ$ $2.36\%$ $2.30\%$ $2.32\%$ $2.34\%$ $2.28\%$ $2.21\%$ $3.77\%$ $3.86\%$ $3.88\%$ $3.82\%$ $3.13\%$ $3.02\%$
Cross Sensor $\circ$ $5.24\%$ $5.20\%$ $5.29\%$ $5.88\%$ $5.34\%$ $5.11\%$ $5.08\%$ $5.11\%$ $5.17\%$ $4.64\%$ $3.91\%$ $3.74\%$
Average $\bullet$ $6.91\%$ $6.85\%$ $6.83\%$ $6.81\%$ $6.55\%$ $6.09\%$ $7.89\%$ $7.99\%$ $8.02\%$ $7.81\%$ $6.56\%$ $6.40\%$
Table 11: The expected calibration error (ECE) of SPVCNN [101] under eight domain-shift scenarios from nuScenes-C and SemanticKITTI-C in the Robo3D benchmark [55]. UnCal, TempS, LogiS, DiriS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, Dirichlet, meta, and our depth-aware scaling calibrations, respectively.
Type nuScenes-C SemanticKITTI-C
UnCal TempS LogiS DiriS MetaC DeptS UnCal TempS LogiS DiriS MetaC DeptS
Clean $\bullet$ $2.57\%$ $2.44\%$ $2.49\%$ $2.54\%$ $2.40\%$ $2.31\%$ $3.46\%$ $2.90\%$ $3.07\%$ $3.41\%$ $2.36\%$ $2.32\%$
Fog $\circ$ $8.53\%$ $8.12\%$ $8.23\%$ $8.54\%$ $7.38\%$ $7.34\%$ $13.06\%$ $12.33\%$ $12.57\%$ $13.23\%$ $11.15\%$ $11.10\%$
Wet Ground $\circ$ $2.80\%$ $2.63\%$ $2.68\%$ $2.72\%$ $2.63\%$ $2.58\%$ $3.52\%$ $3.02\%$ $3.19\%$ $3.49\%$ $2.76\%$ $2.63\%$
Snow $\circ$ $8.49\%$ $7.76\%$ $7.97\%$ $8.35\%$ $6.87\%$ $6.61\%$ $8.50\%$ $7.70\%$ $7.94\%$ $8.41\%$ $6.31\%$ $6.26\%$
Motion Blur $\circ$ $9.18\%$ $8.80\%$ $9.00\%$ $9.33\%$ $8.11\%$ $7.98\%$ $21.01\%$ $19.92\%$ $20.28\%$ $20.41\%$ $17.86\%$ $17.22\%$
Beam Missing $\circ$ $2.88\%$ $2.70\%$ $2.74\%$ $2.79\%$ $2.72\%$ $2.67\%$ $3.01\%$ $2.64\%$ $2.73\%$ $3.04\%$ $2.48\%$ $2.45\%$
Crosstalk $\circ$ $11.76\%$ $11.09\%$ $11.33\%$ $12.01\%$ $9.82\%$ $9.48\%$ $4.66\%$ $4.00\%$ $4.17\%$ $4.49\%$ $3.58\%$ $3.31\%$
Incomplete Echo $\circ$ $2.40\%$ $2.28\%$ $2.33\%$ $2.39\%$ $2.30\%$ $2.24\%$ $3.54\%$ $3.08\%$ $3.24\%$ $3.58\%$ $2.56\%$ $2.52\%$
Cross Sensor $\circ$ $4.80\%$ $4.43\%$ $4.52\%$ $4.57\%$ $4.22\%$ $4.20\%$ $3.27\%$ $2.83\%$ $2.96\%$ $3.36\%$ $2.81\%$ $2.78\%$
Average $\bullet$ $6.36\%$ $5.98\%$ $6.10\%$ $6.34\%$ $5.51\%$ $5.39\%$ $7.57\%$ $6.94\%$ $7.14\%$ $7.50\%$ $6.19\%$ $6.03\%$
Figure 6: The reliability diagrams of randomly sampled model predictions generated by CENet [16] on the validation set of the SemanticKITTI [7] dataset. UnCal, TempS, LogiS, MetaC, and DeptS denote the uncalibrated, temperature, logistic, meta, and our proposed depth-aware scaling calibrations, respectively.
9 Additional Qualitative Result
In this section, we supplement additional qualitative examples to better support the findings and conclusions drawn in the main body of this paper.
9.1 Visualized Calibration Results
We provide additional visualizations to help verify the effectiveness of the proposed model calibration model in enhancing the model’s ability for uncertainty estimation. As can be seen from Fig. 7 and Fig. 8, existing 3D scene understanding models often fail to deliver accurate uncertainty estimates, resulting in potential safety-related issues. Our proposed DeptS is capable of tackling these problems in a holistic manner. After calibration, models can generate more accurate uncertainty estimates, leading to a more reliable 3D scene understanding.
Figure 7: The point-wise expected calibration error (ECE) of existing 3D semantic segmentation models without calibration (UnCal) and with our depth-aware scaling (DeptS). Our approach is capable of delivering accurate uncertainty estimates. The colormap goes from dark to light denotes low and high error rates, respectively.
Figure 8: The point-wise expected calibration error (ECE) of existing 3D semantic segmentation models without calibration (UnCal) and with our depth-aware scaling (DeptS). Our approach is capable of delivering accurate uncertainty estimates. The colormap goes from dark to light denotes low and high error rates, respectively.
10 Limitation and Discussion
In this section, we elaborate on the limitations and potential negative societal impact of this work.
10.1 Potential Limitations
In this work, we established the first benchmark of 3D scene understanding from an uncertainty estimation viewpoint. We also proposed DeptS to effectively calibrate 3D models, achieving more reliable 3D scene understanding. We foresee the following limitations that could be promising future directions.
Data Dependence. Effective model calibration heavily relies on the quality and diversity of the data used. If the dataset is not representative of real-world scenarios or lacks diversity, the calibrated model may not generalize well across different environments or conditions.
Evaluation Challenges. Assessing the effectiveness of calibration can be challenging, as it requires comprehensive metrics that capture the model’s performance across a broad range of scenarios. Standard evaluation metrics may not fully reflect the improvements in reliability and confidence achieved through calibration. It is enlightening to design new metrics for a more holistic evaluation.
10.2 Potential Societal Impact
3D scene understanding often involves capturing and analyzing detailed spatial data about environments which might include private spaces. Additionally, calibrated models might still inherit biases present in the data or algorithmic design, leading to unfair or discriminatory outcomes in certain scenarios. Addressing these issues requires more than technical solutions; it demands careful consideration of the ethical and societal implications of model deployment.
11 Public Resources Used
In this section, we acknowledge the use of the following public resources, during the course of this work.
11.1 Public Codebase Used
We acknowledge the use of the following public codebase during this work:
11.2 Public Datasets Used
We acknowledge the use of the following public datasets during this work:
11.3 Public Implementations Used
We acknowledge the use of the following implementations during this work: